# Lost in Benchmarks? Rethinking Large Language Model Benchmarking with Item Response Theory

### v1

@@ 7 Conclusion
Limitations
Our work still has some limitations: 1) The current analysis is based on a specific selection of contemporary LLMs and benchmarks. Future studies could extend our analysis to encompass a broader and more diverse range of models and datasets as they emerge. 2) PSN-IRT is designed to process binary (correct/incorrect) outcome data. This restricts its direct applicability to benchmarks that utilize graded scoring Zheng et al. (2023), for which extensions incorporating more generalized psychometric models would be necessary. 3) The item parameters estimated by PSN-IRT are currently dependent on the items being included in the training data, meaning that analyzing entirely new, unseen items typically requires retraining the model. Future iterations could explore the integration of item content features to potentially enable parameter estimation for new items without full retraining.
Appendix A Appendix
A.1 Detailed Statistics
@@
A.2 Case Study
@@
  • Examples for discriminability include a quantitative reasoning problem with a high value, indicating strong differentiating power, and a common-sense inference task with a low value, suggesting limited ability to distinguish between models.
@@
A.3 Ablation Study
$P(X=1\mid\theta,a,b)=\frac{1}{1+e^{-a(\theta-b)}}$ (5)
Additionally, we explore replacing one-hot encodings with learned embeddings and substituting the multi-layer perceptron with a graph neural network to model LLM-item interactions. However, these modifications failed to yield significant improvements, likely due to the inherent simplicity of LLM-item interaction patterns that don’t require complex relational modeling. We will explore improved input representations and model architectures as future work to enhance model robustness.
Model Method Parameter ACC F1 AUC Kendall Average
IRT MLE 1PL 0.7219 0.8035 0.6970 1.0000 0.8056
2PL 0.7172 0.7998 0.6963 0.9091 0.7806
3PL 0.7218 0.8031 0.7057 0.9697 0.8001
4PL 0.7211 0.8034 0.7012 0.9697 0.7989
MCMC 1PL 0.7168 0.8011 0.7268 0.9697 0.8036
2PL 0.7232 0.8130 0.7301 0.9697 0.8090
3PL 0.7172 0.8081 0.7297 1.0000 0.8138
4PL 0.7070 0.7811 0.7278 0.9697 0.7964
VI 1PL 0.7209 0.8024 0.6822 0.9697 0.7938
2PL 0.7198 0.8012 0.6938 1.0000 0.8037
3PL 0.7189 0.8011 0.6956 0.9697 0.7963
4PL 0.7201 0.8015 0.6940 0.9091 0.7812
VIBO 1PL 0.7007 0.7801 0.6901 0.9697 0.7852
2PL 0.7301 0.8048 0.6987 0.8788 0.7781
3PL 0.7172 0.7993 0.7048 0.9091 0.7826
4PL 0.7188 0.8007 0.7055 0.9697 0.7987
Deep-IRT Deep Learning 1PL 0.7969 0.8507 0.8507 0.9697 0.8670
PSN-IRT Deep Learning 1PL 0.7948 0.8497 0.8473 0.9091 0.8502
2PL 0.7560 0.8266 0.8259 0.9697 0.8445
3PL 0.7965 0.8510 0.8510 0.8788 0.8443
4PL 0.7991 0.8520 0.8477 1.0000 0.8747

### v2

@@ Results.
In contrast, other item selection approaches are less effective for differentiating among strong models. For instance, selecting items based on discriminability, results in zero score variance when weak models are excluded, indicating it merely separates broad capability tiers rather than providing granularity among top-performers. Separately, a strategy of clustering items based on their success/failure vectors (Pacchiardi, Cheke, and Hernández-Orallo 2024) also fails to improve ranking correlation, even when it increases score separability.
@@ 7 Conclusion
Appendix A Implementation Details
A.1 Estimation Ability of PSN-IRT
To evaluate the estimation ability of PSN-IRT against baselines, we first construct a unified response matrix by aggregating item-level results across all models and benchmarks. We then split these model-item interactions into training (60%), validation (20%), and test sets (20%). The model is trained on the training set using the Adam optimizer with a learning rate of 0.003 and a weight decay of 0.0001, and a batch size of 512. To ensure robust model selection and prevent overfitting, we employ an early stopping strategy with a patience of 5 epochs, monitored based on the F1 score on the validation set.
For the subsequent benchmark analysis presented in Section 5, where the goal is to obtain the most stable parameter estimates for all items, we train the PSN-IRT model on the entire unified dataset. In this phase, the model is trained for a fixed 30 epochs using the same hyperparameters. All training is conducted on a single NVIDIA A800-80G GPU.
A.2 Ablation Study
This section provides additional implementation details for the ablation study variants mentioned in Section 4.3.
Semantic Embedding Variant.
To explore the utility of item content, we substitute one-hot encoded item identifiers with dense semantic embeddings. These embeddings are generated for the text of each benchmark item using the pre-trained Instructor-XL [^10] https://huggingface.co/hkunlp/instructor-xl. They then serve as the direct input to the item network, replacing the one-hot encoding pathway while the rest of the PSN-IRT architecture remained unchanged.
However, these variants fail to yield significant improvements. We will explore improved input representations and model architectures as future work to enhance model robustness.
Graph Neural Network (GNN) Variant.
To investigate the role of network architecture, we implemented a custom GNN-based encoder. This encoder updates a model’s initial embedding by aggregating information from all interactions in the training graph. For each interaction between a model $s$ and an item $q$ with a binary outcome $rel\in\{0,1\}$, a cognitive degree $\mathbf{l}_{s,q}$ is first computed:
$\mathbf{l}_{s,q}=\text{Linear}_{\text{gate}}([\mathbf{e}_{q},\mathbf{W}_{rel}\cdot rel])$ (5)
where $\mathbf{e}_{q}$ is the item’s initial embedding, $\mathbf{W}_{rel}$ is a trainable weight matrix that projects the scalar response $rel$ into the embedding space, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation. This combined vector is then processed by a single linear layer ($\text{Linear}_{\text{gate}}$) to form the cognitive degree. Subsequently, a cognitive state $\mathbf{c}_{s,q}$ is formed through an element-wise product:
$\mathbf{c}_{s,q}=\mathbf{e}_{q}\odot\mathbf{l}_{s,q}$ (6)
The updated embedding for model $s$, denoted $\mathbf{e}^{\prime}_{s}$, is obtained by performing an attention-weighted aggregation of the cognitive states from all items $\mathcal{N}_{s}$ it has interacted with:
$\mathbf{e}^{\prime}_{s}=\mathbf{e}_{s}+\sum_{q_{i}\in\mathcal{N}_{s}}\alpha_{s,q_{i}}\cdot\mathbf{c}_{s,q_{i}}$ (7)
where the attention weights $\alpha_{s,q_{i}}$ are calculated via a softmax over scores. These scores are derived by applying a linear layer and a LeakyReLU activation to a concatenation of the model’s embedding $\mathbf{e}_{s}$ and each cognitive state $\mathbf{c}_{s,q_{i}}$. A symmetric process is used to update item embeddings.
However, both variants fail to yield performance gains. We leave the exploration of more specialized architectures and feature representations as future work.
Appendix B Additional Analysis
B.1 Performance Comparison across IRT Parmaters
$P(X=1\mid\theta,a,b)=\frac{1}{1+e^{-a(\theta-b)}}$ (8)
Model Method Parameter ACC F1 AUC Kendall’s $\tau$
IRT MLE 1PL 0.7219 0.8035 0.6970 1.0000
2PL 0.7172 0.7998 0.6963 0.9091
3PL 0.7218 0.8031 0.7057 0.9697
4PL 0.7211 0.8034 0.7012 0.9697
MCMC 1PL 0.7168 0.8011 0.7268 0.9697
2PL 0.7232 0.8130 0.7301 0.9697
3PL 0.7172 0.8081 0.7297 1.0000
4PL 0.7070 0.7811 0.7278 0.9697
VI 1PL 0.7209 0.8024 0.6822 0.9697
2PL 0.7198 0.8012 0.6938 1.0000
3PL 0.7189 0.8011 0.6956 0.9697
4PL 0.7201 0.8015 0.6940 0.9091
VIBO 1PL 0.7007 0.7801 0.6901 0.9697
2PL 0.7301 0.8048 0.6987 0.8788
3PL 0.7172 0.7993 0.7048 0.9091
4PL 0.7188 0.8007 0.7055 0.9697
Deep-IRT Deep Learning 1PL 0.7974 0.8516 0.8519 0.9697
PSN-IRT Deep Learning 1PL 0.7948 0.8497 0.8473 0.9091
2PL 0.7560 0.8266 0.8259 0.9697
3PL 0.7965 0.8510 0.8510 0.8788
4PL 0.7998 0.8538 0.8485 1.0000
B.2 Detailed Benchmark Statistics
@@
B.3 Relationship Between Item Difficulty and Discriminability
Our analysis reveals a notable relationship between item difficulty and discriminability. As illustrated by the scatter plots in Figure 4, item discriminability tends to be lower when item difficulty reaches extreme levels—that is, when items are either very easy or very hard. Consequently, datasets that feature a more balanced distribution of difficulty often exhibit higher overall discriminability, enabling them to more effectively distinguish between models of varying capabilities.
This pattern is also reflected in the dataset-level averages presented in Table 5 in the main text. For instance, benchmarks such as MATH and GSM8K, which contain a range of difficulties, achieve high average discriminability. Conversely, datasets like GPQA Diamond and TheoremQA, which are characterized by a concentration of very difficult items, exhibit the lowest average discriminability. This suggests that their capacity for precise model differentiation is reduced, in part, due to this difficulty-discriminability trade-off. These observations underscore the importance of balancing difficulty when designing test items to ensure a benchmark is not only challenging but also highly discriminative.
B.4 Case Study
@@
  • For discriminability, the examples include a quantitative reasoning problem with a high value, indicating strong differentiating power, and a common-sense inference task with a low value, suggesting limited ability to distinguish between models.